MOPS – 3(N-morpholino)propanesulfonic acid (structure below) is a common buffer used in biochemistry and biology. Consider a 0.100 M MOPS/Na+•(MOPS-) buffer solution (200.0 mL). pKa (MOPS) = 7.20 (0.1 M MOPS AND 0.1 M Na+ MOPS-)

a. Calculate the pH of the buffer.

b. Calculate the pH when 1.0 mL of 1.0 M NaOH is added to the buffer.

c. Calculate the pH when 5.0 mL of 1.0 M NaOH is added to the buffer.
Answer #1

Given: Concentration of buffer solution = 0.100 M in 200.0 mL

pKa (MOPS) = 7.20

a) Henderson-Hasselbatch equation to calculate pH of the buffer solution is given by

salt pH 3 pКа + log- [аcid]

0.1] pH 7.20 log 0.1]

pH = 7.20

b) MOPS + NaOH -->  Na+MOPS- + H2O

Buffer Initially contains,

MOPS = 0.100 M x 200 mL = 0.002 moles

Na+MOPS- = 0.100 M x 200 mL = 0.002 moles

Addition of 1.0mL of 1.0M NaOH to the buffer

OH- = 0.001L x 1.0mol/L = 0.001 moles

MOPS decreased by 0.001 moles and Na+MOPS- increased by 0.001 moles

[MOPS] = 0.002 -0.001 = 0.001 moles

[Na+MOPS-] = 0.002 + 0.001 = 0.003 moles

0.003 pH 7.20 log 0.001]

pH = 7.20 + log 3

= 7.20 + 0.477

= 7.677

c) MOPS + NaOH -->  Na+MOPS- + H2O

Buffer Initially contains,

MOPS = 0.100 M x 200 mL = 0.002 moles

Na+MOPS- = 0.100 M x 200 mL = 0.002 moles

Addition of 5.0mL of 1.0M NaOH to the buffer

OH- = 0.005L x 1.0mol/L = 0.005 moles

MOPS decreased by 0.005 moles and Na+MOPS- increased by 0.005 moles

[MOPS] = 0.002 -0.005 = -0.003 moles

[Na+MOPS-] = 0.002 + 0.005 = 0.007 moles

0.007 pH 7.20 log, -0.003

pH = 7.20 - log 2.33

= 7.20 - 0.367

= 6.833
Answer #2

Formula is right. He converted wrong. 200ml is 0.2L. Just adjust accordingly.

answered by: JOIDSF